# Continuum models describing probabilistic motion of tagged agents in exclusion processes

### v1

@@ II.4 Numerical methods
$\frac{\partial C^{(s)}}{\partial t}+\dfrac{\partial\mathcal{J}_{x}^{(s)}}{\partial x}=\mathcal{S}^{(s)},$ (12)
@@
$\frac{\partial P^{(s)}}{\partial t}+\dfrac{\partial\mathcal{J}^{(s)}}{\partial x}=0,$ (14)
@@
$\mathcal{J}_{x}^{(s)}$ $=-D^{(s)}\left(1-T\right)\dfrac{\partial P^{(s)}}{\partial x}-D^{(s)}P^{(s)}\dfrac{\partial T}{\partial x}+v_{x}^{(s)}P^{(s)}\left(1-T\right).$ (15)
@@
$\frac{\partial C}{\partial t}$ $=D\frac{\partial^{2}C}{\partial x^{2}}-v_{x}\frac{\partial}{\partial x}\left[C(1-C)\right],$ (16)
$\frac{\partial P}{\partial t}$ $=D\left[(1-C)\frac{\partial^{2}P}{\partial x^{2}}+P\frac{\partial^{2}C}{\partial x^{2}}\right]-v_{x}\frac{\partial}{\partial x}\left[P(1-C)\right],$ (17)
with no-flux boundary conditions for $C(x,t)$ and $P(x,t)$, at $x=\pm W/2$. We solve this system of coupled PDEs using the method of lines by uniformly discretising the $x$ coordinate with a constant step size $h$ and approximating the spatial derivatives in the PDE models using finite differences. For the second-order diffusive derivatives in Equations (16)–(17), we use the standard central difference approximation. For the first-order advective derivatives, we use an upwind scheme, which approximates the derivative of a function at $x$ in terms of the values of the function at $x$ and at xhx-h for $v_{x}>0$ or values of the function at $x$ and $x+h$ for $v_{x}<0$. We consider the initial condition
$C(x,0)$ $=\left\{\begin{array}[]{ll}C_{0},&|x|\leq a,\\ 0,&\textrm{otherwise},\end{array}\right$. (20)
$P(x,0)$ $=\left\{\begin{array}[]{ll}\dfrac{1}{h},&x=x_{0},\\ 0,&\textrm{otherwise},\end{array}\right$. (23)
@@
$F(x,t)=\int_{-W/2}^{x}P(x^{\prime},t)\,\textrm{d}x^{\prime}.$ (24)
@@
All simulations use a lattice of width $W=300$, height $H=100$, and initial density $C_{0}=1$ in the region $|x|\leq 20$, i.e. Equation (20) with $a=20$. In each ABM simulation, we place a set of 10 tagged agents at each of three initial locations: $x_{0}=-18$ (close to the left-most leading edge); $x_{0}=0$ (in the centre of the population); and $x_{0}=18$ (close to the right-most leading edge). The initial $y$-coordinate of the tagged agents is chosen randomly (without replacement) from $y=0,1,\ldots,H$. This design of the placement of agents replicates the experimental design taken by Cai et al. [26] shown in Figure 1(c) where five cells are chosen along each vertical transect with constant horizontal position, $x$. To generate averaged data we perform $5000$ independent realisations of the stochastic ABM, and calculate the average column occupancies, and the proportion of each group of tagged agents in each column, over all identically prepared realisations. Matlab software to reproduce the results in this study is publicly available at [33]. All results presented here use Matlab R2022b. We encourage readers can use this software directly to replicate the results presented in this study, or to adapt the software and explore different scenarios, such as working with different initial conditions, parameter values, simulation durations.
III Results
In this section we compare stochastic ABM simulations to numerical solutions of the PDE models to illustrate how the PDE models for $P(x,t)$ can be used to predict the distribution of tagged agent locations from stochastic simulations. We systematically explore the behaviour of the models with and without direction bias, and with and without agent proliferation.
III.1 Unbiased motility and no proliferation
Figure 2: Comparison of discrete and continuum models for a motile population without bias or proliferation. (a) Agent density $C(x,t)$ at $t=300$. (b) Distribution at of the location at $t=300$ of tagged agents initially located near the left-hand leading edge ($x_{0}=-18$, blue), in the centre of the population ($x_{0}=0$, red) and near the right-hand leading edge ($x_{0}=18$, yellow). Vertical dashed lines show the initial location of tagged agents. (c) Median and 90% PrI of tagged agent locations as a function of time: ABM results are shown as thick solid curve (median) and shaded band (90% PrI); PDE results are shown as thin solid curve (median) and dashed curves (90% PrI). Discrete parameter values $M=1$ and $\rho_{x}=Q=0$ corresponding to $D=0.25$ and $v=\lambda=0$ for a simulation with $\Delta=\tau=1$.
III.2 Biased motility and no proliferation
When there is a directional bias in movement in the positive $x$ direction with $\rho_{x}>0$, the macroscopic agent density becomes skewed to the right (Figure 3a), as is well known [27]. Again, we see that averaged simulation data for the density matches the solution of the mean-field density, including the skew in the density profile. The distribution of tagged agents initially near the right-most leading edge drifts in the positive $x$-direction due to the motility bias, and diffuses over time (Figure 3b, yellow). Tagged agents initially in the centre $x=0$ and at the left-most leading edge drift more slowly and do not spread out as rapidly due to crowding effects (Figure 3b, red, blue). Again, the PDE gives a good approximation to the distribution of tagged agent locations in the ABM (Figure 3b-c).
Figure 3: Comparison of discrete and continuum models for a motile population with bias and without proliferation. (a) Agent density $C(x,t)$ at $t=300$. (b) Distribution of the location at $t=300$ of tagged agents initially located near the left-hand leading edge ($x_{0}=-18$, blue), in the centre of the population ($x_{0}=0$, red) and near the right-hand leading edge ($x_{0}=18$, yellow). Vertical dashed lines show the initial location of tagged agents. (c) Median and 90% PrI of tagged agent locations as a function of time: ABM results are shown as thick solid curve (median) and shaded band (90% PrI); PDE results are shown as thin solid curve (median) and dashed curves (90% PrI). Discrete parameter values $M=1$, $\rho_{x}=0.2$ and $Q=0$ corresponding to $D=0.25$, $v=0.1$ and $\lambda=0$ for a simulation with $\Delta=\tau=1$.
III.3 Unbiased motility and proliferation
Including proliferation in the ABM, with no directional bias, means that the total population size increases over time and the mean-field PDE model is the well-known Fisher-Kolmogorov model [34, 35, 36]. Therefore, after a sufficient duration of time the macroscopic agent density eventually approaches a travelling wave profile with $C(x,t)\to 1^{-}$ (representing a fully occupied lattice) behind the wave front and $C(x,t)\to 0^{+}$ ahead of the leading edge (Figure 4a). The distributions of tagged agents initially close to either leading edge start to spread out at the beginning of the simulation, and eventually become skewed away from the centre (Figure 4b, blue and yellow), as occurred in the case without proliferation in Figure 2. However, these agents that are initially at the leading-ledge of the population eventually become relatively immobile and fixed in place once the lattice in their local neighbourhood becomes fully occupied. As a result, there is little further change in their distributions after approximately $t=150$ (Figure 4c). The tagged agents initially located in the centre of the population near $x=0$ are unable to move at all and so retain their initial Dirac delta-like distribution for the whole simulation (Figure 4b-c, red). Again, the solution of the PDE model for $P(x,t)$ captures these trends and dynamics reasonably accurately.
Figure 4: Comparison of discrete and continuum models for a motile population without biased and with proliferation. (a) Agent density $C(x,t)$ at $t=300$. (b) Distribution of the location at $t=300$ of tagged agents initially located near the left-hand leading edge ($x_{0}=-18$, blue), in the centre of the population ($x_{0}=0$, red) and near the right-hand leading edge ($x_{0}=18$, yellow). Vertical dashed lines show the initial location of tagged agents. (c) Median and 90% PrI of tagged agent locations as a function of time: ABM results are shown as thick solid curve (median) and shaded band (90% PrI); PDE results are shown as thin solid curve (median) and dashed curves (90% PrI). Discrete parameter values $M=1$, $\rho_{x}=0$ and $Q=0.025$ corresponding to $D=0.25$, $v=0$ and $\lambda=0.025$ for a simulation with $\Delta=\tau=1$.
III.4 Biased motility and proliferation
In simulations with both directional bias and proliferation, the macroscopic agent density develops an asymmetric profile, with the travelling wave at the right-most leading edge moving faster than that at the left-most leading edge (Figure 5a). Tagged agent distributions display varying degrees of drift and diffusion depending on their initial location within the population (Figure 5b). Tagged agents in the centre are unable to move and are fixed in place almost immediately due to high local density (Figure 5b-c, red). Those tagged agents initially near the left-hand edge spread out until around $t=150$ when they become immobile (Figure 5b-c, blue), whereas the tagged agents initially near the right-most leading edge are able to move more rapidly and their distribution is continuing to spread out at the end of the simulation at $t=300$ (Figure 5b-c, yellow). This is because some of these tagged agents are able to remain ahead of the wave front in a region where macroscopic density remains sufficiently small. Again the PDE provides a good approximation to the ABM results.
Figure 5: Comparison of discrete and continuum models for a motile population with biased and proliferation. (a) Agent density $C(x,t)$ at $t=300$. (b) Distribution of the location at $t=300$ of tagged agents initially located near the left-hand leading edge ($x_{0}=-18$, blue), in the centre of the population ($x_{0}=0$, red) and near the right-hand leading edge ($x_{0}=18$, yellow). Vertical dashed lines show the initial location of tagged agents. (c) Median and 90% PrI of tagged agent locations as a function of time: ABM results are shown as thick solid curve (median) and shaded band (90% PrI); PDE results are shown as thin solid curve (median) and dashed curves (90% PrI). Discrete parameter values $M=1$, $\rho_{x}=0.2$ and $Q=0.025$ corresponding to $D=0.25$, $v=0.1$ and $\lambda=0.025$ for a simulation with $\Delta=\tau=1$.
III.5 Accuracy of the approximation for tagged agents
We explored how the accuracy of the PDE model compared to the method proposed by Simpson, Landman and Hughes [27] for calculating the mean and variance of tagged agent locations. This method, which we refer to as the SLH approximation, relies on calculating the rate of change of the mean and variance for an agent that is positioned at the mean location, and thus ignores the effects of uncertainty in the agent location. In all cases investigated, the PDE model and the SLH approximation both predict the mean agent location very accurately (Supplementary Figures S1a,c,e,g). In cases without proliferation, the PDE model predicts the variance as well or better than the SLH approximation (Supplementary Figure S1b,d). When there is proliferation, the results are more mixed. When there is no bias, the SLH predicts the variance of agents initially near a leading edge more accurately than the PDE model (Supplementary Figure S1f), but the reverse is true when there is bias (Supplementary Figure S1h). The PDE model tends to overestimate the variance in tagged locations in all cases, while the SLH approximation underestimates it in some cases and overestimates in others.
Whilst we have calculated the SLH approximation for all cases investigated here for completeness, it should be noted that Simpson et al. [27] did not consider proliferation, and only examined agents initially located near the right-most leading edge of the population. Also the SLH approximation provides no information about tagged agent locations beyond their mean and variance, whereas our model provides access to the full distribution via its probability density function.
To illustrate this, we explored a test case in which all lattice sites were initially occupied with probability $0.5$ and there is no proliferation or motility bias. This means that the macroscopic agent density is uniformly constant at $C(x,t)=0.5$ and the PDE for $P(x,t)$ in Equation (17) reduces to the linear diffusion equation with diffusivity $0.5D$. This PDE for $P(x,t)$ would accurately describe a situation where 50% of the tagged agent’s attempted movement events are aborted. However, simulations of the ABM show that the tagged agent’s distribution spreads out more slowly than this PDE predicts (Supplementary Figure S2), indicating that more than 50% of moves are aborted. This is due to the positive correlation between neighbouring lattice sites, which means that, conditional on the tagged agent being at site $(i,j)$, the occupancy probabilities of sites $(i\pm 1,j)$ and $(i,j\pm 1)$ are slightly greater than $0.5$.
IV Conclusion and Future Work
In this work, we have investigated lattice-based random walk models of individual-level motility and proliferation mechanisms in the context of an exclusion process framework, where each lattice site can be occupied by no more than a single agent [21]. Stochastic models with exclusion are often used to represent cell biology experiments where crowding effects can be very pronounced [1, 2, 3, 4, 7, 8, 9, 10, 11, 13, 14, 15]. Motivated by experimental observations in Figure 1, we used a stochastic ABM to describe the spatial evolution of the population-level density profile, as well as considering the motion of a small number of tagged agents within the broader population. We explored the existing continuum-limit description of the macroscopic density $C(x,y,t)\in[0,1]$ alongside a new continuum-limit description of the probability density function for the location of tagged agents $P(x,y,t)\geq 0$. The derivation of the continuum limit was extended to the situation where we have $S\in\mathbb{N}^{+}$ distinct subpopulations on the lattice so that the dependent variables in the PDE models are $C^{(s)}(x,y,t)\in[0,1]$ and $P^{(s)}(x,y,t)\geq 0$ for $s=1,2,3,\ldots,S$. Repeated stochastic simulation data showed that the continuum-limit model for $C$ and $P$ provide a good match data obtained from the computationally expensive stochastic simulations. In particular, we showed that numerical estimates of $P$ provide a probabilistic interpretation of the motion of tagged agents since the continuum-limit model can be used to predict both the expected location of the tagged agents, as well as predicting the variability in their location. Numerical tests confirmed that the solution of the continuum-limit model for $P$ provides a reasonable match to data from the stochastic ABM under a range of conditions including unbiased and biased motility, both in the presence and absence of agent proliferation.
The PDE models we derived for the density $C^{(s)}$ involve a source term whenever $Q^{(s)}>0$ for $s=1,2,\ldots,S$. The corresponding PDE for the probability density function for tagged agents $P^{(s)}$ does not involve any source term, even when the simulation involves agent proliferation with $Q^{(s)}>0$. This property is a reflection of the fact that our model tracks the location of agents that are tagged at the beginning of the experiment, without consideration of any offspring they may produce. Our approach could be generalised to track the original tagged agents, as well as any daughter agents and associated lineages they give rise to. This would lead to a PDE model for $P^{(s)}$ that involves a source term to account for the fact that tagged agents themselves can be involved in proliferation events, similar to mathematical models of lineage tracing [37].
Another approach for generalising our mathematical models is to recast the stochastic lattice-based random walk model into a stochastic lattice-free random walk [38, 39, 40]. The main advantage of working in a lattice-free framework is that the motility direction is continuous rather than being discrete, however this additional flexibility comes at the cost of additional mathematical complexity, which makes deriving appropriate mean-field PDE descriptions more difficult [38, 39, 40].
Regardless of whether we work with a lattice-based or lattice-free framework, deriving appropriate continuum limits for the population-level density and individual-level tagged agent trajectory properties provides an opportunity to perform parameter inference using simultaneous observations of the agent density, $C^{(s)}$, and the location of tagged agents, $P^{(s)}$ for $s=1,2,\ldots,S$. Having simple computational tools that enable us to understand how $C^{(s)}$ and $P^{(s)}$ vary with the model parameters provides a key ingredient for either Bayesian or frequentist inference [41, 42] that makes full use of all available (combined) data. Having a continuum-limit PDE description is highly advantageous for parameter inference as: (1) when coupled with a suitable observation noise model [30, 41], it provides access to a likelihood function that can be used for optimisation or sampling-based methods; and (2) it is more computationally efficient than generating repeated stochastic ABM simulations, which can be prohibitively expensive for performing parameter inference, especially if the parameter space is high-dimensional.
Acknowledgements This work is partly supported by the Australian Research Council (DP230100025, CE230100001) and the Marsden Fund (24-UOC-020).
Supplementary Material

### v2

@@ 4 Conclusion and Future Work
$\frac{\partial C^{(s)}}{\partial t}+\dfrac{\partial\mathcal{J}_{x}^{(s)}}{\partial x}=\mathcal{S}^{(s)},$ (S1)
@@
$\frac{\partial P^{(s)}}{\partial t}+\dfrac{\partial\mathcal{J}^{(s)}}{\partial x}=0,$ (S3)
@@
$\mathcal{J}_{x}^{(s)}$ $=-D^{(s)}\left(1-T\right)\dfrac{\partial P^{(s)}}{\partial x}-D^{(s)}P^{(s)}\dfrac{\partial T}{\partial x}+v_{x}^{(s)}P^{(s)}\left(1-T\right).$ (S4)
@@
$\frac{\partial C}{\partial t}$ $=D\frac{\partial^{2}C}{\partial x^{2}}-v_{x}\frac{\partial}{\partial x}\left[C(1-C)\right],$ (S5)
$\frac{\partial P}{\partial t}$ $=D\left[(1-C)\frac{\partial^{2}P}{\partial x^{2}}+P\frac{\partial^{2}C}{\partial x^{2}}\right]-v_{x}\frac{\partial}{\partial x}\left[P(1-C)\right],$ (S6)
with no-flux boundary conditions for $C(x,t)$ and $P(x,t)$, at $x=\pm W/2$. We solve this system of coupled PDEs using the method of lines by uniformly discretising the $x$ coordinate with a constant step size $h$ and approximating the spatial derivatives in the PDE models using finite differences. For the second-order diffusive derivatives in Equations (S5)–(S6), we use the standard central difference approximation. For the first-order advective derivatives, we use an upwind scheme, which approximates the derivative of a function at $x$ in terms of the values of the function at $x$ and at xhx-h for $v_{x}>0$ or values of the function at $x$ and $x+h$ for $v_{x}<0$. We consider the initial condition
$C(x,0)$ $=\left\{\begin{array}[]{ll}C_{0},&|x|\leq a,\\ 0,&\textrm{otherwise},\end{array}\right$. (S9)
$P(x,0)$ $=\left\{\begin{array}[]{ll}\dfrac{1}{h},&x=x_{0},\\ 0,&\textrm{otherwise},\end{array}\right$. (S12)
@@
$F(x,t)=\int_{-W/2}^{x}P(x^{\prime},t)\,\textrm{d}x^{\prime}.$ (S13)
@@
All simulations use a lattice of width $W=300$, height $H=100$, and initial density $C_{0}=1$ in the region $|x|\leq 20$, i.e. Equation (S9) with $a=20$. In each ABM simulation, we place a set of 10 tagged agents at each of three initial locations: $x_{0}=-18$ (close to the left-most leading edge); $x_{0}=0$ (in the centre of the population); and $x_{0}=18$ (close to the right-most leading edge). The initial $y$-coordinate of the tagged agents is chosen randomly (without replacement) from $y=0,1,\ldots,H$. This design of the placement of agents replicates the experimental design taken by Cai et al. [26] shown in Figure 1(c) where five cells are chosen along each vertical transect with constant horizontal position, $x$. To generate averaged data we perform $5000$ independent realisations of the stochastic ABM, and calculate the average column occupancies, and the proportion of each group of tagged agents in each column, over all identically prepared realisations.
Matlab software to reproduce the results in this study is publicly available at [35]. All analyses were run in Matlab R2022b.